Three years ago, Jeff was twice as old as Mike is now. In nine years, Jeff will be four years older than Mike is now. What are their current ages?

Answer by stanbon(75887)

Three years ago, Jeff was twice as old as Mike is now. In nine years, Jeff will be four years older than Mike is now. What are their current ages?
---------------------------------------------------------
Let jeff's age now be "j".
Let mike's age now be "m".
-------------------------------
Equations:
j-3 = 2m
j+9= m+4
---------------
Rearrange:
j = 2m+3
Substitute:
2m+3 + 9 = m+4
m = -8 (this is an unrealistic answer; please check your problem statement)
==========================================
Cheers,
Stan H.